OpenStudy (anonymous):

find an equation of the line containing the given pair of points {1/6,-1/3} and {5/6,3} y=?

OpenStudy (anonymous):

y=4.97x-1.14

OpenStudy (anonymous):

(y-3) = (x- 5/6) * (-1/3 -3)/(1/6- 5/6)

jimthompson5910 (jim_thompson5910):

Slope \[m=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{3-(-\frac{1}{3})}{\frac{5}{6}-\frac{1}{6}}\] \[m=\frac{3+\frac{1}{3}}{\frac{5}{6}-\frac{1}{6}}\] \[m=\frac{\frac{9}{3}+\frac{1}{3}}{\frac{5}{6}-\frac{1}{6}}\] \[m=\frac{\frac{9+1}{3}}{\frac{5-1}{6}}\] \[m=\frac{\frac{10}{3}}{\frac{4}{6}}\] \[m=(\frac{10}{3})(\frac{6}{4})\] \[m=\frac{60}{12}\] \[m=5\] --------------- \[y-y_1=m(x-x_1)\] \[y-3=5(x-\frac{5}{6})\] \[y-3=5x-\frac{25}{6}\] \[y=5x-\frac{25}{6}+3\] \[y=5x-\frac{7}{6}\]